# A Mapper algorithm with implicit intervals and its optimization

### v1

@@ 3.2 The Mapper graph mode
Here, the $Q_{i}^{*}$ and $Q_{i}^{**}$ are the largest and the second largest elements of $i$-th row of $Q$, respectively. ${i}^{*}$ and ${i}^{**}$ are their corresponding indices. For $i=1,\ldots,n,$, the mode of the $i$-th row of the hidden assignment matrix is then determined as follows (see Appendix A for the theoretical derivation).
@@ 3.3 Loss function
The probability matrix $Q$ is derived from a Gaussian mixture model fitted to the projected data. Therefore, optimizing the Mapper graph is equivalent to optimizing the parameter $\boldsymbol{\theta}$ of GMM. However, the maximum likelihood estimation of parameter $\boldsymbol{\theta}$ only accounts for the distributions of the projected data, neglecting the topological information of the Mapper graph. To construct an appropriate Mapper graph, it is essential to consider both the likelihood of the projected data and the topology of data simultaneously. We design a loss function that incorporates both the likelihood of projected data and topological information of the Mapper graph. The log-likelihood of the projected data is given by
@@
We adopt a similar formula to Oulhaj et al. (2024) to measure the topological information loss. The topological information of a Mapper graph can be encoded by the extended persistence diagram (Carrière and Oudot, 2018). A significant advantage of the extended persistence diagram is its ability to capture the branches of a Mapper graph. The branches of a Mapper graph are crucial in downstream data analysis. We denote the extended persistence diagram as $D$, where $D=\{(b_{i},d_{i})\mid i=1,\dots,M\}$ is a set contains $M$ points. We use a function $Pers:G\mapsto D$ to represent the process of computing extended persistence diagram for a given Mapper graph.
@@
To represent the topological information on the extended persistence diagram, a function that maps the diagram to a real number is defined, denoted as $l:D\mapsto\mathbb{R}$. This function is called the persistence specific function, which measures the topological information captured in the extended persistence diagram. There are various functions available for representing overall topological information, such as persistence landscapes (Bubenik, 2015) or computing the bottleneck distance from a target extended persistence diagram (Bauer et al., 2024). In this work, we adopt the averaged persistence which measure the averaged persistence time of each point as the persistence-specific loss,
$l(D)=\frac{1}{M}\sum_{i=1}^{M}|d_{i}-b_{i}|$
This value offers a reasonable summary of the extended persistence diagram, where points with short persistence times are considered noise, and those with long persistence times are regarded as signals. An effective diagram should predominantly feature signal points.
The extended persistence diagram $D$ is derived from a Mapper graph $G$, $G$ is generated by the Mapper function of the hidden assignment matrix $H$. For notation simplicity, we denote
@@
Assume there are $M$ features concerning the Mapper graph mode. The total loss function of the GMM soft Mapper can be defined as a weighted average of the negative log-likelihood of the filtered data and the topological loss concerning $H$
$Loss(\boldsymbol{\theta}|\mathbb{X}_{n},\mathbb{Y}_{n})$ $=-\lambda_{1}\frac{\log L(\mathbb{Y}_{n}|\boldsymbol{\theta})}{n}-\lambda_{2}l(H_{mode})$
Here, the first term is the averaged log-likelihood, representing the information carried out by the filter data. By averaging, we ensure that the log-likelihood is comparable across datasets of varying sample sizes. The denominator in the second term is used to avoid encouraging spurious structures in the Mapper graph. The parameters $\lambda_{1}$ and $\lambda_{2}$ control the relative weights of the likelihood of the projected data and the topological loss of the Mapper graph.
@@ 3.4 Stochastic gradient descent parameter estimation
In this study, we utilize stochastic gradient descent algorithm (SGD) (Bottou, 2010) to optimize the loss function defined above. Our goal is to find parameters $\boldsymbol{\theta}$ for a GMM that minimize the total loss function. The optimization process aims to improve the topological structure of a Mapper graph while ensuring high likelihood of the filtered data. Once the optimal parameters are determined, we can either sample from the GMM to create a soft Mapper graph or directly compute the Mapper graph mode. Notably, our methods does not require the derivation of an explicit gradient expression. This simplification is facilitated by automatic differentiation frameworks such as PyTorch (Paszke et al., 2019) or TensorFlow (Abadi et al., 2016), which effectively implement SGD.
@@
Instead of updating $\pi_{i}$ directly, this approach updates $\xi_{i}$. This transformation guarantees that $\pi_{i}$ meets the constraints at each step. Similarly, to satisfy the constraint of standard variance $\sigma_{i}>0$, we take the transformation $\log\sigma_{i}=\xi^{\prime}_{i}$. Then we update $\xi^{\prime}_{i}$ instead of $\sigma_{i}$ at each step. The detailed SGD algorithm for the GMM soft Mapper model is presented in Algorithm 1.
Input: Learning rate $\gamma$, steps $N$, weights $\lambda_{1},\lambda_{2}$, number of intervals $K$,
@@ 4 Synthetic datasets
In this section, we compare our proposed algorithm to the standard Mapper algorithm on synthetic datasets to demonstrate its effectiveness. The output graphs indicate that our proposed method perform comparable or better than the standard Mapper algorithm concerning the topological structures. In order to illustrate the uncertainty in the soft Mapper graph, we also output a few samples from the optimized distribution (see Appendix B). Training loss of the GMM soft Mapper while optimizing prameter for each dataset is given in the Appendix D.
@@ 4.1 Uniform dataset
We start with two uniformly sampled datasets. $1000$ data points are sampled from two separate circles and data sampled from two overlapping circles, respectively. Figures 2 (a) and (e) provide a visualization of these two datasets. In both datasets, the filter function is the $x$-axis coordinates. We employ the DBSCAN clustering algorithm with parameters set to $\epsilon=0.6$ and $minPts=5$. For the disjoint circles dataset, we set the number of intervals to $K=6$ and for the intersecting circles dataset, we set $K=5$. Full details of the parameters are listed in Table 1. The Mapper graph modes, both with and without optimization, are depicted in Figures 2 (c), (d), and (g), (h). To illustrate the inherent uncertainty in the soft Mapper, we also present eight samples from the optimized distribution in the supplemental materials. The results demonstrate that the Mapper graph modes effectively approximate the topological structures of these datasets.
@@
As the underlying data structure for these two small datasets is relatively simple, both the standard Mapper graph and the optimized Mapper graph mode can easily reveal the underlying structure. The topological structure of the data can even be well captured by the Mapper graph mode with an appropriate choice of the GMM initial values without optimization.
@@
Figure 2: Comparison of the standard Mapper algorithm and our proposed algorithm on a two disjoint circles dataset and a two intersecting circles dataset. (a, e). A visualization of the datasets. Coloured lines represents the intervals produced by each algorithm. (b, f). Output graphs of the standard Mapper algorithm with $n=6,p=0.33$ for the two disjoint circles and $n=5,p=0.2$ for the two intersecting circles. (c, g). The output Mapper graph mode of our proposed method without optimization. (d, h). The output Mapper graph mode of our proposed method with optimization.
@@
We make slight modifications to the datasets above, resulting in two disjoint and two intersecting circles with different radius, each containing $1000$ points. The filter function remains the $x$-axis coordinates. For the unequal-sized disjoint circles, DBSCAN parameters are set to $\epsilon=0.3$ and $minPts=5$; for the unequal-sized intersecting circles, parameters are $\epsilon=0.2$ and $minPts=5$. The standard Mapper struggles to capture the correct topological structure due to its fixed interval setting, whereas our method can adjust intervals based on data distribution, facilitating the accurate capture of the data’s underlying shape. In the case of the unequal-sized intersecting circles dataset, the Mapper graph mode fails to correctly capture the data structure without optimization. However, with optimization, the intervals are slightly adjusted, leading to accurate results.
@@
Figure 3: Comparison of the standard Mapper algorithm and our proposed algorithm on a two unequal-sized disjoint circles dataset and a two unequal-sized intersecting circles dataset. (a, e). A visualization of the datasets. Coloured lines represents the intervals produced by each algorithm. (b, f). Output graphs of the standard Mapper algorithm with $n=6,p=0.1$ for the disjoint circles and $n=5,p=0.2$ for the intersecting circles. (c, g). The output Mapper graph mode of our proposed method without optimization. (d, h). The output Mapper graph mode of our proposed method with optimization.
@@ 4.2 Noisy dataset
In this section, we assess the robustness of our proposed algorithm by introducing noises into the intersecting circles dataset. The noise is drawn from a Gaussian distribution and is added to each point’s coordinates. Initially, we introduce a small amount of noise, characterized by a Gaussian distribution with mean of zero and standard deviation of $0.1$. The dataset is shown in Figure 4 (a). With this setting for noises, both the standard Mapper and our Mapper graph mode can accurately represent the correct topological structure. The results are shown in Figure 4 (b)-(d).
Subsequently, we increase the noise level, increasing the standard deviation up to $0.3$. At this noise level, the standard Mapper algorithm is unable to output meaningful structures, as illustrated in Figure 4 (f). Conversely, after optimization, our Mapper graph mode still can capture the two primary loops, as shown in Figure 4 (h), though with some additional branches and isolated points. These results highlight the efficacy of flexible interval partitioning in uncovering the data’s inherent structure, even in the presence of significant noise. Parameter optimization plays a crucial role in accurately determining the intervals, which boosts the algorithm’s performance.
Figure 4: Comparison of the standard Mapper algorithm and our proposed algorithm on a two intersecting circles dataset with noises. (a, e). A visualization of the datasets. Coloured lines represents the intervals produced by each algorithm. (b, f). Output graphs of the standard Mapper algorithm with $n=6,p=0.4$ for the disjoint circles and $n=6,p=0.4$ for the intersecting circles. (c, g). The output Mapper graph mode of our proposed method without optimization. (d, h). The output Mapper graph mode of our proposed method with optimization.
@@
We finally test our proposed algorithm on a 3D human dataset from Oulhaj et al. (2024). In this example, the number of intervals is set to $K=8$ and the DBSCAN clustering with parameters $\epsilon=0.6,minPts=5$ is used. The filter function is the mean value of distance between each sample and others. The Mapper graph mode, initialized with an appropriate choice for the GMM parammeters, captures the topological structure better than the standard Mapper graph as shown in Figure 5. After optimization, the Mapper graph mode shows a more concise structure. These results indicates that the proposed soft Mapper algorithm performs better than the standard Mapper concerning the visualization of topological structures, and the optimization process successfully further optimize the topological structures.
Figure 5: Comparison of the standard Mapper algorithm and our proposed algorithm on a 3D human dataset. (a) A visualization of the 3D human dataset. (b) The projected data and resulted intervals of each algorithm. (c) The output graph of the standard Mapper algorithm, the number of intervals is $8$ and the overlap rate is $0.1$. (d) The Mapper graph mode without optimization. (e) The Mapper graph mode with optimization.

### v2

@@ 3.2 The Mapper graph mode
Here, the $Q_{i}^{*}$ and $Q_{i}^{**}$ are the largest and the second largest elements of $i$-th row of $Q$, respectively. ${i}^{*}$ and ${i}^{**}$ are their corresponding indices. For $i=1,\ldots,n$, the mode of the $i$-th row of the hidden assignment matrix is then determined as follows (see Appendix A for the theoretical derivation).
@@ 3.3 Loss function
The probability matrix $Q$ is derived from a GMM fitted to the projected data. Therefore, optimizing the Mapper graph is equivalent to optimizing the parameter $\boldsymbol{\theta}$ of the GMM. However, the maximum likelihood estimation of parameter $\boldsymbol{\theta}$ only accounts for the distributions of the projected data, neglecting the topological information of the Mapper graph. To construct an appropriate Mapper graph, it is essential to consider both the likelihood of the projected data and the topology of data simultaneously. We design a loss function that incorporates both the likelihood of projected data and topological information of the Mapper graph. The log-likelihood of the projected data is given by
@@
We adopt a similar formula to Oulhaj et al. (2024) to measure the topological information loss. The topological information of a Mapper graph can be encoded by the extended persistence diagram (Carrière and Oudot, 2018). A significant advantage of the extended persistence diagram is its ability to capture the branches of a Mapper graph. The branches of a Mapper graph are crucial in downstream data analysis. We denote the extended persistence diagram as $D$, where $D=\{(b_{i},d_{i})\mid i=1,\dots,M\}$ is a multiset containing $M$ points. We use a function $Pers:G\mapsto D$ to represent the process of computing the extended persistence diagram for a given Mapper graph.
@@
To represent the topological information on the extended persistence diagram, a function that maps the diagram to a real number is defined, denoted as $l:D\mapsto\mathbb{R}$. This function is called the persistence specific function, which measures the topological information captured in the extended persistence diagram. There are various functions available for representing overall topological information, such as persistence landscapes (Bubenik, 2015) or computing the bottleneck distance from a target extended persistence diagram (Bauer et al., 2024). Not all points on a persistence diagram are meaningful, some points may generated through noises in the dataset. Typically, points with short persistence are considered noises, while those with long persistence are regarded as meaningful signals. An effective diagram should primarily consist of signal points, reflecting the significant topological structures of the dataset. The Mapper graph intermediates between the data and the persistent homology, so that a good Mapper graph has not just (relatively) small noise but in fact few features attributable to noise. To get a more robust Mapper graph, in this work, we adopt the averaged persistence which measures the averaged persistence time of each point as the persistence-specific loss, assume there are $M$ features concerning the Mapper graph mode, the topological loss is defined as:
$l(D)=\frac{1}{M}\sum_{i=1}^{M}|d_{i}-b_{i}|.$
This value offers a reasonable summary of the extended persistence diagram. An effective diagram should predominantly feature signal points. The detonator in this term is used to avoid encouraging spurious structures in the Mapper graph.
The extended persistence diagram $D$ is derived from a Mapper graph $G$, while $G$ is generated by the Mapper function of the hidden assignment matrix $H$. For notation simplicity, we denote
@@
The total loss function of the GMM soft Mapper can be defined as a weighted average of the negative log-likelihood of the filtered data and the topological loss concerning $H$
$Loss(\boldsymbol{\theta}|\mathbb{X}_{n},\mathbb{Y}_{n})$ $=-\lambda_{1}\frac{\log L(\mathbb{Y}_{n}|\boldsymbol{\theta})}{n}-\lambda_{2}l(H_{mode}).$
Here, the first term is the averaged log-likelihood, representing the information carried out by the filter data. By averaging, we ensure that the log-likelihood is comparable across datasets of varying sample sizes. The parameters $\lambda_{1}$ and $\lambda_{2}$ control the relative weights of the likelihood of the projected data and the topological loss of the Mapper graph.
@@ 3.4 Stochastic gradient descent parameter estimation
In this study, we utilize stochastic gradient descent algorithm (SGD) (Bottou, 2010) to optimize the loss function defined above. Our goal is to find parameters $\boldsymbol{\theta}$ for a GMM that minimize the total loss function. The optimization process aims to improve the topological structure of a Mapper graph while ensuring high likelihood of the filtered data. Once the optimal parameters are determined, we can either sample from the GMM to create a soft Mapper graph or directly compute the Mapper graph mode. Notably, our method does not require the derivation of an explicit gradient expression. This simplification is facilitated by automatic differentiation frameworks such as PyTorch (Paszke et al., 2019) or TensorFlow (Abadi et al., 2016), which effectively implement SGD.
@@
Instead of updating $\pi_{i}$ directly, this approach updates $\xi_{i}$. This transformation guarantees that $\pi_{i}$ meets the constraints at each step. Similarly, to satisfy the constraint of standard variance $\sigma_{i}>0$, we take the transformation $\log\sigma_{i}=\xi^{\prime}_{i}$. Then we update $\xi^{\prime}_{i}$ instead of $\sigma_{i}$ at each step. The detailed SGD algorithm for the GMM soft Mapper model is presented in Algorithm 1. In this work, we implement the proposed algorithm in Python version 3.10. The extended persistence diagram and bottleneck distance is computed by Python package GUDHI version 3.8.0, the stochastic gradient descent is implemented through pytorch 1.13.
Input: Learning rate $\gamma$, steps $N$, weights $\lambda_{1},\lambda_{2}$, number of implicit intervals $K$,
@@ 4 Synthetic datasets
In this section, we compare our proposed algorithm to the standard Mapper algorithm and D-Mapper on synthetic datasets to demonstrate its effectiveness. The output graphs indicate that our proposed method performs comparable or better than the standard Mapper and D-Mapper algorithm concerning the topological structures. To quantitatively compare these algorithms, we calculate the Silhouette Coefficient ($SC$) (Han et al., 2012) and the adjusted $SC$ ($SC_{adj}$) (Tao and Ge, 2024) for each dataset in Section 4.4, see Table 2. In order to illustrate the uncertainty in the soft Mapper graph, we also output a few samples from the optimized distribution (see Appendix B, figure 7 to figure 13), and provide some summary statistics concerning some summary metrics of graphs in Table 3 in Appendix B. Training loss of the GMM soft Mapper while optimizing prameter for each dataset is given in Appendix E.
@@ 4.1 Uniform dataset
We start with two uniformly sampled datasets. $1000$ data points are sampled from two separate circles and data sampled from two overlapping circles, respectively. Figures 2 (a) and (f) provide a visualization of these two datasets and (implicit) intervals produced by each algorithm. To visualize the implicit intervals, we take the minimum point of each cluster as the starting point of the interval and the maximum point of each cluster as the ending point of the interval. In both datasets, the filter function is the $x$-axis coordinates. For the disjoint circles dataset, we set the number of (implicit) intervals to $K=6$ and apply the DBSCAN clustering algorithm with parameters $\epsilon=0.3$ and $minPts=5$. For the intersecting circles dataset, we use $K=5$ with the same clustering algorithm with parameters $\epsilon=0.2$ and $minPts=5$. Full details of the parameters are listed in Table 1. The Mapper graph modes, both with and without optimization, are depicted in Figures 2 (d), (i), and (e), (j). The output graphs demonstrate that all the Mapper graphs effectively approximate the topological structures of these datasets.
@@
As the underlying data structure for these two small datasets is relatively simple, all these three algorithms can easily reveal the underlying structure. The topological structure of the data can even be well captured by the Mapper graph mode with an appropriate choice of the GMM initial values without optimization.
@@
Figure 2: Comparison of the standard Mapper algorithm, D-Mapper algorithm and our proposed algorithm on a two disjoint circles dataset and a two intersecting circles dataset. (a, f). A visualization of the datasets. Coloured lines represents the (implicit) intervals produced by each algorithm. (b, g). The output graphs of the standard Mapper algorithm with $K=6,p=0.33$ for the two disjoint circles and $K=5,p=0.2$ for the two intersecting circles. (c, h). The output graphs of D-Mapper algorithm with $\alpha=0.08$ for the disjoint circles and $\alpha=0.31$ for the intersecting circles. (d, i). The output Mapper graph mode of our proposed method without optimization. (e, j). The output Mapper graph mode of our proposed method with optimization.
@@
We make slight modifications to the datasets above, resulting in two disjoint and two intersecting circles with different radii, each containing $1000$ points. The filter function remains the $x$-axis coordinates. For the unequal-sized disjoint circles, DBSCAN parameters are set to $\epsilon=0.35$ and $minPts=5$, for the unequal-sized intersecting circles, parameters are $\epsilon=0.2$ and $minPts=5$. The results of the three algorithms are shown in Figure 3. The standard Mapper struggles to capture the correct topological structure due to its fixed interval setting, whereas the D-Mapper and our method can adjust intervals based on data distribution, facilitating the accurate capture of the data’s underlying shape. In the case of the unequal-sized intersecting circles dataset, the Mapper graph mode fails to correctly capture the data structure without optimization. However, after optimization, the implicit intervals (groups) are slightly adjusted, leading to more accurate results.
@@
Figure 3: Comparison of the standard Mapper algorithm, D-Mapper algorithm and our proposed algorithm on a two unequal-sized disjoint circles dataset and a two unequal-sized intersecting circles dataset. (a, f). A visualization of the datasets. Coloured lines represents the (implicit) intervals produced by each algorithm. (b, g). The output graphs of the standard Mapper algorithm with $K=6,p=0.3$ for the disjoint circles and $K=5,p=0.2$ for the intersecting circles. (c, h). The output graphs of D-Mapper algorithm with $\alpha=0.06$ for the disjoint circles and $\alpha=0.05$ for the intersecting circles. (d, i). The output Mapper graph mode of our proposed method without optimization. (e, j). The output Mapper graph mode of our proposed method with optimization.
@@ 4.2 Noisy dataset
In this section, we assess the robustness of our proposed algorithm by introducing noises into the intersecting circles dataset. The noise is drawn from a Gaussian distribution and is added to each point’s coordinates. Initially, we introduce a small amount of noise, characterized by a Gaussian distribution with mean of zero and standard deviation of $0.1$. The dataset is shown in Figure 4 (a). With this setting for noises, both the standard Mapper, D-Mapper and our Mapper graph mode can accurately represent the correct topological structure. The results are shown in Figure 4 (b)-(e).
Subsequently, we increase the noise level, increasing the standard deviation up to $0.3$. The dataset is shown in Figure 4 (f). At this noise level, the standard Mapper algorithm is unable to output meaningful structures, as illustrated in Figure 4 (g). The D-Mapper algorithm can captuare the main two loops of the data, while still has some extra branches and an isolated node, as shown in Figure 4 (h). After optimization, our Mapper graph mode still can capture the two primary loops, as shown in Figure 4 (j), though with some additional branches. These results highlight the efficacy of flexible implicit interval partitioning in uncovering the data’s inherent structure, even in the presence of significant noise. The parameter optimization plays a crucial role in accurately determining the implicit intervals, which boosts the algorithm’s performance.
Figure 4: Comparison of the standard Mapper algorithm, D-Mapper algorithm and our proposed algorithm on a two intersecting circles dataset with noises. (a, f). A visualization of the datasets. Coloured lines represents the (implicit) intervals produced by each algorithm. (b, g). The output graphs of the standard Mapper algorithm with $K=6,p=0.4$ for the small noise dataset and $K=6,p=0.4$ for the big noise dataset. (c, h). The output graphs of D-Mapper algorithm with $\alpha=0.025$ for the small noise dataset and $\alpha=0.07$ for the big noise dataset. (d, i). The output Mapper graph mode of our proposed method without optimization. (e, j). The output Mapper graph mode of our proposed method with optimization.
@@
We finally test our proposed algorithm on a 3D human dataset from Oulhaj et al. (2024). In this example, the number of (implicit) intervals is set to $K=8$ and the DBSCAN clustering with parameters $\epsilon=0.1,minPts=5$ is used. The filter function is the mean value of distance between each sample and others. The standard Mapper outputs an asymmetrical skeleton, while both the D-Mapper and the Mapper graph mode without optimization produce a symmetrical skeleton. After optimization, the Mapper graph mode outputs a more concise structure (a symmetric skeleton with less nodes) than the standard Mapper graph and D-Mapper graph as shown in Figure 5. These results indicates that the proposed soft Mapper algorithm performs better than the standard Mapper concerning the visualization of topological structures, and the optimization process successfully further optimize the topological structures.
Figure 5: Comparison of the standard Mapper algorithm, D-Mapper and our proposed algorithm on a 3D human dataset. (a) A visualization of the 3D human dataset. (b) The projected data and resulted (implicit) intervals of each algorithm. (c) The output graph of the standard Mapper algorithm, the number of intervals is $8$ and the overlap rate is $0.1$. (d) The output graph of D-Mapper algorithm, the parameter $\alpha=0.071$. (e) The Mapper graph mode without optimization. (f) The Mapper graph mode with optimization.
4.4 Quantitatively summary of synthetic datasets
Through a series of synthetic dataset experiments, we find that the optimized Mapper graph mode can capture dataset shapes more accurately in general. It should be noted that the quantitative evaluation of the Mapper graphs is a challenging problem, as the topological information of a given data is often not available (Bui et al., 2020; Carrière et al., 2018). As in most of the literature, here we compute the $SC$, this metric accounts for the clustering performance only (Bui et al., 2020; Chalapathi et al., 2021; Dłotko, 2019). The $SC_{adj}$ accounts for both the clustering performance as well as the topological information, to some extent (Tao and Ge, 2024).
As shown in Table 2, in all cases but the two intersecting dataset and the 3D human dataset, the D-Mapper performs comparably or better than the than the standard Mapper and the Mapper graph mode performs best in terms of metric $SC_{adj}$. For the two intersecting circles dataset, the $SC_{adj}$ of the D-Mapper is lower than the standard Mapper graph. This is mainly caused by the lower $TSR$ value of the D-Mapper. This dataset indicates that the D-Mapper algorithm may be unstable when the number of intervals $K$ ($=5$) is small for some cases, and when $K$ is set to a large number ($K=8$), denoted as D-Mapper* in Table 2, a larger $TSR$ is obtained. The graph of D-Mapper* and more explanations can be found in Appendix C. For the 3D human dataset, the $SC_{adj}$ of the Mapper graph mode is a little bit lower than that of the standard Mapper and D-Mapper, but as the output graphs indicates (see Figure 5), the Mapper graph mode has more condensed structures.